# Understanding Codebase like a Professional! Human–AI Collaboration for Code Comprehension

### v2

@@ Abstract.
To address this, we propose learning from the perspective of a unique role, code auditors, whose work often requires them to quickly familiarize themselves with new code projects on weekly or even daily basis. To achieve this, we recruited and interviewed 8 code auditing practitioners to understand how they master codebase understanding. We identified several design opportunities for an LLM-based codebase understanding system: supporting cognitive alignment through automated codebase information extraction, decomposition, and representation, as well as reducing manual effort and conversational distraction through interaction design.
To validate these four design opportunities, we designed a system prototype, CodeMap, that provides dynamic information extraction and representation aligned with the human cognitive flow and enables interactive switching among hierarchical codebase visualizations. To evaluate the usefulness of our system, we conducted a user study with nine experienced developers and six novice developers. Our results demonstrate that CodeMap improved users’ perceived intuitiveness, ease of use, and usefulness in supporting code comprehension, while reducing their reliance on reading and interpreting LLM responses by 79% and increasing map usage time by 90% compared with the static visualization analysis tool. It also enhances novice developers’ perceived understanding and reduces their unpurposeful exploration. The insights derived from our interviews and the design of CodeMap can inspire future LLM-based research on code comprehension.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
A decent understanding on the codebase is the foundation for developers to work productively. A research found that on average developers spend around 58% of their time on program comprehension activities (Xia et al., 2017). Especially, as software evolves, codebases are growing larger and becoming increasingly complex, gain understanding become more critical to properly perform further software development and maintenance operations. Codebase comprehension remain challenging even in the era of large language models (LLMs). For example, Reefman et al. (Reefman, 2024) found that LLM tools cannot provide the business logic necessary to understand a codebase. Similarly, Alebachew et al. (Berhanu Alebachew, 2025) pointed out that LLM tools also fail to offer the adaptive, dynamic, and interactive strategies that developers use in real-world settings.
To systematically aid developers (regardless experienced and novice level) in codebase-level understanding, we first observed that a special role, i.e., domain practitioners in code auditing (as known as code auditors), whose job is to frequently onboard new codebases (i.e., weekly or daily) to examine the codebase quality, such as inspecting potential vulnerabilities and write reports (Contributors, 2025)—possess valuable tacit knowledge that experienced and novice developers (i.e., onboarding monthly or quarterly) may lack. This tacit knowledge, with best practice, is valuable and could benefit the developers’ activities related to codebase understanding. Motivated by the above, we aim to answer the following key question:
@@
Through thematic analysis, we identified key patterns in how code auditors understand codebases: their understanding flows from global to local levels, and along this flow, they sequentially seek information about the project overview, codebase structure and business logic, and local functions and variables. We also observed how they used ChatGPT to support this process and the challenges they encountered while doing so. Based on these findings, we derived key design opportunities: LLM systems should facilitate information extraction, decomposition, and representation that align with this hierarchical understanding flow, and support users’ interaction within it. Finally, we implemented a prototype based on identified design opportunities, and user validation with both experienced and novice developers showed that CodeMap enhanced overall user experience and increased feature engagement.
@@ 2.1. Why Code Auditor?
Code Auditing is the thorough analysis of source code to identify security vulnerabilities and deviations from best practices. It can be traced back to 1976, when Michael Fagan (Fagan, 1976) formalized a six-step code inspection workflow that involves a thorough process of reading and understanding the codebase. Specifically, auditors must often familiarize themselves with an unfamiliar project both globally (holistic enough) and line by line (detailed enough) on a weekly or even daily basis. Through this process, auditors must develop a deep, structured grasp of the codebase because their vulnerability‑detection task demands it.
@@
Our observation in this difference between code auditors and developers has made us realized that auditors may have tacit knowledge that general developers could learn in performing codebase understanding process.
@@ 2.2. Research Questions
To answer our key research question described in the Introduction, we propose four detailed research questions (RQs) to further guide our study. By learning from code auditors, we are interested in:
  • RQ1: a) How do auditors cognitively navigate and understand an unfamiliar codebase? b) What are their information retrieval patterns during the process?
  • RQ2: a) What types of assistance do they often seek from LLMs (e.g., ChatGPT) during the process? b) What challenges do they encounter? c) What suggestions do they provide?
Moreover, we are also interested in whether the findings derived could a set of actionable insights to further inspire the design of future LLM-powered Codebase Understanding System, particularly, we aim to answer:
  • RQ3: How does our expert-inspired system affect both experienced and novice developers’ perceived usefulness, intuitiveness, ease of use, and others?
@@ 3. Methodology Overview
Figure 1. Methodology overview.
To answer our questions, as shown in Figure 1, we employed a four-stage method from formative interview to prototype development and validation, which is commonly used in human-computer interaction (HCI) research (Helander, 2014; Group, 2019).
@@
  • Stage 2: Design opportunities development. To support the future development of LLM-powered Codebase Understanding System, we translated our identified professional auditing strategies into four design opportunities through which the LLM system could provide cognitive support in information extraction, representation, reasoning, and user interaction.
  • Stage 3: Prototype design. To operationalize our design opportunities, we designed and developed an LLM-powered web application prototype, CodeMap. It incorporates two key features: (1) dynamically extracting, decompose, representing information aligned with human’s hierarchical understanding flow; and (2) allowing users to interact and switch between different abstraction layers, updating the visualization as their understanding evolves.
  • Stage 4: Prototype validation. To validate the effectiveness of the prototype, we conducted an evaluation with 9 experienced developers and 6 novice developers as the real users. The final results from this stage answer RQ3 and RQ4 by collecting participants’ subjective perceptions through post-study questionnaires and analyzing their verbal feedback during system use, across both experienced and novice developers.
@@ 4.1. Method
ID Occupation
@@
Exp. (Y)
Audit
Exp. (Y)
Software
No. of Audited
Projects
Onboarding
Frequency
P1 SC Auditor Daily 7 3 200–250 Daily to Weekly
P2 SC Auditor Daily 2 2 100-150 Weekly
P3 Engineer Weekly 3 1.5 100–150 Weekly
P4 SC Auditor Daily 2.5 2.5 100–150 Weekly
P5 SC Auditor Daily 2 2 50–100 Weekly
P6 CS Student Daily 3 1 0–50 Monthly
P7 SC Auditor Daily 4 1 0–50 Weekly to Monthly
P8 SC Auditor Monthly 3 0.5 0–50 Monthly
Notes: SC Auditor = ”Smart Contract Auditor”; CS Student = ”Computer Science Master
Student” Coding Platforms (e.g., auditwizard.io) Assistance Tools (e.g., VSCode
plugins) – Auditing Tools (e.g., slither, echidna, certora)
During the interview, we asked participants to reflect upon and share their most notable code auditing practices and experiences of using LLMs to aid code auditing. In addition to narrating their experiences, we requested participants to demonstrate their code auditing process via screen sharing when feasible. We also explored the challenges they encountered with LLMs and their expectations concerning potential AI assistance in the process. All interviews were conducted virtually, audio-recorded, and transcribed for further subsequent analysis.
Following Richards et al.’s collaborative approach to qualitative analysis (Richards and Hemphill, 2018), the interviewer and another qualitative researcher jointly analyzed the data. First, we independently performed open coding on three representative transcripts. Next, we met to compare codes, resolve disagreements, and agree on an initial codebook. We then piloted and refined this codebook on additional data. Finally, one coder applied the finalized codebook to the remaining transcripts. The full codebook and detailed results are available on our website (Authors, [n. d.]). We present findings below.
@@
Understanding Flow and Information Seeking ChatGPT Assistance (), Challenges (p) & User Suggestions (ȷ) ChatGPT’s Assistance Example Quotes
Global Level Understanding. Project Overview. Gathering high-level information (e.g., name of key files, project background and auditing requirements) from README files and documentation. ”After obtaining the project, I will read its written content, including white paper and README. I will not read very carefully. I only need to know what each specific contract is for.” P1 Support surface level high-level understanding;
Evaluate project complexity
p Lack of automated guidance for codebase understanding (need prompting)
ȷ System can provide pre-defined domain prompts
”For example, [to assess] how many days we can save and how much money we can charge for this project. Previously, it might have taken two or three days to assess, but now we may only need one or two minutes to get the job done [via ChatGPT].” P6
Global Level Understanding. Code structure and logic: business logic flow and function-call flow. Focus on the main components, their relationships, and the key files, functions, and variables. E.g., show how files are connected through imports and what is imported; examine inheritance, the entry point, and the data flow; and describe how classes and functions interact. ”You have a general basic [structure] framework… Some companies will draw a diagram, such as architecture diagram.” P6 Understand which component is using this variable;
p Lack of automated guidance for codebase understanding (need prompting)
p Lack of codebase information decomposition (need manual decomposition)
p Irrelevant code analysis drives redundant effort
ȷ System can provide codebase structure visualization
ȷ System can provide pre-defined domain prompts
@@
”Eespecially when I want to understand the relationships between multiple files or multiple contracts, it probably won’t be able to help me. Especially when ChatGPT has more input data, the feedback it provides is rather broad.” P7
Local Level Understanding. Locate components, files, variables, and specific contents users are interested: Important variables, key variable’ dependency order, summarize important functions, specify important information. ”For example, there might be an issue [at a certain point], and I’ll make a note indicating what kind of issue it is… Then, after finishing looking at this [overall structure], I will come back and focus on this marked part. That’s roughly how it goes.” P8 Extract information from single document
p Lack of support for cross file variable relationships
p Lack of codebase information decomposition (need manual decomposition)
p Decide which part the attention should go to
”For example, at the beginning, I want it to tell me the specific content of the entire contract on a large scale. I will specifically tell it to tell me what variables are inside, that is, some important state variables, and help me summarize the important functions of those public functions. I will probably specify for it to summarize certain content”. P7
Details Understanding. Clarify understanding of code snippet: Whether the specific content is similar to the users or not Explain code snippet
Confirm user’s understanding
p Clarification question often break understanding flow
”After giving ChatGPT a piece of code, [I may have] many questions. After you ask question a, when you then ask question b, there is a process interruption. If you ask question b, it will combine the answer to question a to give you the answer to question b.” P5
@@ 4.2. Results
RQ1: How do auditors cognitively navigate and understand an unfamiliar codebase? What information are they seeking during this process?
@@
RQ2: What types of assistance did they seek from LLMs during the process? What challenges did they encounter, and what suggestions did they provide?
Throughout this process, we find that auditors frequently relied on conversational LLMs such as ChatGPT for summarizing or explaining code fragments. However, they reported three recurring challenges: (1) the need for repetitive manual prompting to retrieve hierarchical information, (2) the cognitive burden of manually decomposing large projects into promptable segments, and (3) irrelevant or overly general answers that failed to align with their reasoning focus. As one participant noted, “I can ask the model to explain a function, but it never tells me how this connects to the rest of the system.
Our results highlight that while current LLM tools can localize detailed information, they lack mechanisms to represent structural relationships and support auditors’ layered reasoning flow. Code auditors envisioned tools that could automatically organize LLM—generated information according to their cognitive workflow—providing visual overviews, structured decompositions, and explanations that maintain cross-level coherence.
@@ 5. Stage 2: Design Opportunities for LLM-powered Codebase Understanding System
Considering onboarding process remain challenging nowadays, we hypothesize that developers may not acquire these skills in understanding a new codebase effectively. Due to the practical needs, great complexity of the task, and high learning costs, developers may rely on a LLM system that can provide systematic guidance in the understanding process. Below, we summarize the insights on design opportunities for developing an LLM-powered system to effectively support codebase understanding (also in Figure 2)..
@@
Figure 2. Human-LLM co-understanding codebase. LLM-powered codebase understanding system could provide effective assistance in human cognitive understanding by offering improved information extraction, representation, and decomposition.
@@
Design Opportunity 1: LLM-powered Codebase Understanding System should support information extraction that aligns with the human understanding flow. Instead of directly displaying all information from the codebase, our interviews reveal that users need to manually gather specific information at different levels: global level—project background and structural information; local level—key files and naming information; and detailed level—information clarification. This suggests that an LLM-powered Codebase Understanding System should be designed to support a structured three-level chain of comprehension, reducing users’ manual efforts to seek information across these levels.
@@
Design Opportunity 2: LLM-powered Codebase Understanding System should support information decomposition to align with the human mental model, deepening understanding within each abstraction layer. Our interview findings show that users need to perform information decomposition—breaking the structure into different functional components—so they can focus on specific parts for deeper understanding. This process is often manual and requires expertise. Therefore, an LLM-powered Codebase Understanding System should assist with information decomposition to reduce the manual effort required from users.
@@
Design Opportunity 3: LLM-powered Codebase Understanding System should support information representation connecting across multiple abstraction layers. Information in a codebase is inherently complex, spanning multiple files, functional components, and detailed function implementations. Grasping these relationships requires substantial human cognitive effort, particularly when reasoning about inter-module dependencies. Therefore, an LLM-powered Codebase Understanding System should be able to provide structural representations at multiple abstraction layers, enabling easier and more efficient comprehension.
@@
Design Opportunity 4: LLM-powered Codebase Understanding System should decreasing manual efforts in formulating LLM queries. Our interview findings show that users need both expertise and skill to formulate effective LLM queries—especially to locate the specific information they wish to focus on and to translate their analytical goals into language that the LLM can interpret to provide useful explanations. However, this query formulation process often distracts users from their original flow of understanding. Therefore, an LLM-powered Codebase Understanding System should support additional detailed queries that operate outside the main understanding flow, allowing users to flexibly transition between the system’s structured flow of understanding and their own evolving reasoning structure. Moreover, novice developers may not possess the same analytical skills or domain expertise as professionals. To reduce their cognitive and formulation burden, an LLM-powered Codebase Understanding System could provide predefined information structures or commonly requested details that are automatically displayed to users, allowing them to access relevant insights without the need to manually formulate queries.

### v3

@@ Abstract.
To address this, we propose learning from the perspective of a unique role, code auditors, whose work often requires them to quickly familiarize themselves with new code projects on a weekly or even daily basis. To achieve this, we recruited and interviewed 8 code auditing practitioners to understand how they master codebase understanding. We identified four design opportunities for an LLM-based codebase understanding system: supporting cognitive alignment through automated codebase information extraction, decomposition, and representation, as well as reducing manual effort and conversational distraction through interaction design.
To validate these four design opportunities, we designed a system prototype, CodeMap, that provides dynamic information extraction and representation aligned with the human cognitive flow and enables interactive switching among hierarchical codebase visualizations. To evaluate the usefulness of our system, we conducted a user study with nine experienced developers and six novice developers. Our results demonstrate that CodeMap significantly improved users’ perceived intuitiveness, ease of use, and usefulness in supporting code comprehension, while reducing their reliance on reading and interpreting LLM responses by 79% and increasing map usage time by 90% compared to the static visualization analysis tool. It also enhances novice developers’ perceived understanding and reduces their unpurposeful exploration. The insights derived from our interviews and the design of CodeMap can inspire future LLM-based research on code comprehension, such as onboarding support systems.
Code Auditing, Codebase Understanding, Large Language Models, Understanding Chain, Code Visualization
journalyear: 2026copyright: ccconference: 34th IEEE/ACM International Conference on Program Comprehension; April 12–13, 2026; Rio de Janeiro, Brazilbooktitle: 34th IEEE/ACM International Conference on Program Comprehension (ICPC ’26), April 12–13, 2026, Rio de Janeiro, Brazildoi: 10.1145/3794763.3794822isbn: 979-8-4007-2482-4/2026/04ccs: Human-centered computing Visualization systems and tools
@@ 1. Introduction
A decent understanding on the codebase is the foundation for developers to work productively. A study found that on average developers spend around 58% of their time on program comprehension activities (Xia et al., 2017). Especially, as software evolves, codebases are growing larger and becoming increasingly complex, gaining understanding becomes more critical to properly perform further software development and maintenance operations. Codebase comprehension remains challenging even in the era of large language models (LLMs). For example, Reefman et al. (Reefman, 2024) found that LLM tools cannot provide the business logic necessary to understand a codebase. Similarly, Alebachew et al. (Alebachew, 2025) pointed out that LLM tools also fail to offer the adaptive, dynamic, and interactive strategies that developers use in real-world settings.
To systematically aid developers (regardless of experience level) in codebase-level understanding, we first observed that a special role, i.e., domain practitioners in code auditing (also known as code auditors), whose job is to frequently onboard new codebases (i.e., weekly or daily) to examine the codebase quality, such as inspecting potential vulnerabilities and write reports (Contributors, 2025)—possess valuable tacit knowledge that experienced and novice developers (i.e., onboarding monthly or quarterly) may lack. This tacit knowledge, with best practice, is valuable and could benefit the developers’ activities related to codebase understanding. Motivated by the above, we aim to answer the following key question:
@@
Through thematic analysis, we identified key patterns in how code auditors understand codebases: their understanding flows from global to local levels, and along this flow, they sequentially seek information about the project overview, codebase structure and business logic, and local functions and variables. We also observed how they used LLMs (e.g., ChatGPT)  [^1] This interview study was conducted in 2024, when ChatGPT was among the most advanced tools available for AI-assisted code reading. More recent tools such as Cursor and Trae may offer somewhat different capabilities, but the core strategies identified here remain the same. to support this process and the challenges they encountered while doing so. Based on these findings, we derived key design opportunities: LLM systems should facilitate information extraction, decomposition, and representation that align with this hierarchical understanding flow, and support users’ interaction within it. Finally, we implemented a prototype based on identified design opportunities, and user validation with both experienced and novice developers showed that CodeMap enhanced overall user experience and increased feature engagement.
Figure 1. Methodology overview.
@@ 2.1. Why Code Auditor?
Code Auditing is the thorough analysis of source code to identify security vulnerabilities and deviations from best practices. It can be traced back to 1976, when Michael Fagan introduced the ”Fagan Inspection” (Fagan, 1976), which formalized a six‑step code inspection workflow that reduced defects up to 90%. It often contains thorough codebase understanding process. Specifically, auditors must often familiarize themselves with an unfamiliar project both globally (holistic enough) and line by line (detailed enough) on a weekly or even daily basis. Through this process, auditors must develop a deep, structured grasp of the codebase because their vulnerability‑detection task demands it.
@@
Our observation in this difference between code auditors and developers has made us realize that auditors may have tacit knowledge that general developers could learn in performing codebase understanding process.
@@ 2.2. Research Questions
To answer our key research question described in Section 1, we propose four detailed research questions (RQs) to further guide our study. By learning from code auditors, we are interested in:
  • RQ1: How do auditors cognitively navigate and understand an unfamiliar codebase? What information are they seeking during the process?
  • RQ2: What types of assistance do they seek from LLMs (e.g., ChatGPT) during the process? What challenges do they encounter? What suggestions do they have?
Moreover, we are also interested in whether the findings derived could yield a set of actionable insights to further inspire the design of future LLM-powered Codebase Understanding System, particularly, we aim to answer:
  • RQ3: How does our expert-inspired system affect experienced and novice developers’ perceived intuitiveness, ease of use, and others?
@@ 3. Methodology Overview
To answer our questions, as shown in Figure 1, we employed a four-stage method from formative interview to prototype development and validation, which is commonly used in HCI research (Helander, 2014; Group, 2019).
@@
  • Stage 2: Design opportunities. To support the future development of LLM-powered Codebase Understanding System, we translated our identified professional auditing strategies into four design opportunities through which the LLM system could provide cognitive support in information extraction, decomposition, representation, and user interaction.
  • Stage 3: Prototype design. To operationalize our design opportunities, we designed and developed an LLM-powered web application prototype, CodeMap. It incorporates two key features: (1) dynamically extracting, decomposing, representing information aligned with human’s hierarchical understanding flow; and (2) allowing users to interact and switch between different abstraction layers, updating the visualization as their understanding evolves.
  • Stage 4: Prototype validation. To validate the effectiveness of the prototype, we conducted an evaluation with 9 experienced developers and 6 novice developers. Results from this stage address RQ3 and RQ4, drawing on participants’ subjective perceptions captured through post-study questionnaires and verbal feedback elicited during system use.
@@ 4.1. Method
ID Occupation Tools Used
Onboarding
Frequency
No. of Audited
Projects
@@
Exp. (Years)
Auditing
Exp. (Years)
P1 Smart Contract Auditor Coding Platforms; VSCode Plugin; Auditing Tools Daily to Weekly 200–250 Daily 7 3
P2 Smart Contract Auditor VSCode Plugin; Auditing Tools Weekly 100–150 Daily 2 2
P3 Engineer VSCode Plugin; Auditing Tools Weekly 100–150 Weekly 3 1.5
P4 Smart Contract Auditor VSCode Plugin; Auditing Tools Weekly 100–150 Daily 2.5 2.5
P5 Smart Contract Auditor VSCode Plugin Weekly 50–100 Daily 2 2
P6 Computer Science Master Student Auditing Tools Monthly 0–50 Daily 3 1
P7 Smart Contract Auditor VSCode Plugin; Auditing Tools Weekly to Monthly 0–50 Daily 4 1
P8 Smart Contract Auditor VSCode Plugin Monthly 0–50 Monthly 3 0.5
Notes: Coding Platforms (e.g., auditwizard.io); Auditing Tools (e.g., slither, echidna, certora)
During the interview, we asked participants to reflect upon and share their most notable code auditing practices and experiences of using LLMs to aid code auditing. In addition to narrating their experiences, we requested participants to demonstrate their code auditing process via screen sharing when feasible. We also explored the challenges they encountered with LLMs and their expectations concerning potential AI assistance in the process. All interviews were conducted virtually, audio-recorded, and transcribed for further analysis.
Following Richards et al.’s collaborative approach to qualitative analysis (Richards and Hemphill, 2018), the interviewer and another qualitative researcher jointly analyzed the data  [^3] Prior work has also explored LLM-assisted approaches to collaborative qualitative analysis (Gao et al., 2024b). We chose a fully manual coding process to ensure analytical rigor and maintain close engagement with the data.. First, we independently performed open coding on three representative transcripts. Next, we met to compare codes, resolve disagreements, and agree on an initial codebook. We then piloted and refined this codebook on additional data. Finally, one coder applied the finalized codebook to the remaining transcripts. The full codebook and detailed results are available on our website. We present findings below.
@@
Understanding Flow and Information Seeking ChatGPT Assistance ($\checkmark$), Challenges ($\times$) & User Suggestions ($\star$) Example Quotes (Slightly modified for readability)
Global Level Understanding. Auditors gather project overview information, including key file names, project background, and auditing requirements, from README files and documentation. ”After obtaining the project, I will read its written content, including white paper and README. I will not read very carefully. I only need to know what each specific contract is for.” P1 $\checkmark$ Assist in high-level explanation of the codebase;
$\checkmark$ Assist in evaluating codebase complexity
$\times$ Lack of automated guidance for codebase understanding (requires manual prompting)
$\star$ LLM tools should provide pre-defined domain-specific prompts
”For example, [to assess] how many days we can save and how much money we can charge for this project. Previously, it might have taken two or three days to assess, but now we may only need one or two minutes to get the job done [via LLMs].” P6
Global Level Understanding. Auditors understand business logic flow and function-call flow, seeking main components, relationships, key files, functions, and variables. For example, they examine how files are connected through imports, what inheritance relationships exist, where the entry point is, how data flows through the codebase, and how classes and functions interact. ”You have a general basic [structure] framework… Some companies will draw a diagram, such as architecture diagram.” P6 $\checkmark$ Assist in understanding which components use a given variable;
$\times$ Lack of automated guidance for codebase understanding (requires manual prompting)
$\times$ Lack of automated information decomposition (requires manual decomposition)
$\times$ Lack of filtering for irrelevant code, leading to redundant analysis effort
$\star$ LLM tools should provide pre-defined domain prompts
$\star$ LLM tools should provide codebase structure visualization
@@
”Especially when I want to understand the relationships between multiple files or multiple contracts, it probably won’t be able to help me. Especially when I input more data to ChatGPT, the response is very broad.” P7
Local Level Understanding. Auditors locate components, files, variables, and specific contents of interest, such as important variables and their dependency order, key functions and their call relationships. ”There might be an issue [at a certain point], and I’ll make a note saying what kind of issue it is… Then, after finishing looking at this [overall structure], I will come back and focus on this marked part.” P8 $\checkmark$ Assist in extracting key information from individual files
$\times$ Lack of support for cross-file variable relationships
$\times$ Lack of automated information decomposition (requires manual decomposition)
$\star$ LLM tools should help users identify which parts to focus on
”At the beginning, I want it to explain to me the specific content of the codebase. I will tell it to explain to me the important state variables, and help me summarize the important functions. I will probably also ask it to summarize certain content”. P7
Detailed Level Understanding. Auditors seek clarification of specific code snippets, verifying whether the code aligns with their current understanding. $\checkmark$ Assist in explaining code snippets
$\checkmark$ Assist in confirming users’ understanding
$\times$ Lack of continuous understanding flow—clarification questions often interrupt the ongoing reasoning process
”After giving ChatGPT a code snippet, [I may have] many questions. After you ask question a, when you then ask question b, there is a process interruption. If you ask question b, it will combine the answer to question a to give you the answer to question b.” P5
@@ 4.2. Results
RQ1: How do auditors cognitively navigate and understand an unfamiliar codebase? What information are they seeking during the process?
@@
RQ2: What types of assistance do they seek from LLMs (e.g., ChatGPT) during the process? What challenges did they encounter? What suggestions do they have?
Throughout this process, we find that auditors frequently relied on conversational LLMs such as ChatGPT for summarizing or explaining code fragments. However, they reported three recurring challenges: (1) the need for repetitive manual prompting to retrieve and understand hierarchical structure, (2) the cognitive burden of manually decomposing large projects into promptable segments, and (3) irrelevant or overly general answers that failed to align with their reasoning focus. As one participant noted, “I can ask the model to explain a function, but it never tells me how this connects to the rest of the codebase.
Our results highlight that while current LLM tools can explain detailed information, they lack mechanisms to represent structural relationships and support auditors’ layered cognitive flow. Code auditors envisioned tools that could automatically organize LLM-generated information according to their cognitive workflow—providing visual overviews, structured decompositions, and explanations that maintain cross-level coherence.
@@ 5. Stage 2: Design Opportunities for LLM-powered Codebase Understanding System
Since the onboarding process remains challenging, developers may not easily acquire effective codebase understanding skills. Given the practical demands, task complexity, and high learning costs, developers may benefit from an LLM system that provides systematic guidance throughout the understanding process. Below, we summarize the design opportunities derived from our formative study for developing an LLM-powered system to effectively support codebase understanding (also in Figure 2).
@@
Figure 2. Human-LLM co-understanding of codebases. An LLM-powered codebase understanding system can support human cognitive processes by automating information extraction, decomposition, and structural representation.
@@
Design Opportunity 1: LLM-powered Codebase Understanding System should support information extraction aligned with the human understanding flow. Instead of directly displaying all information from the codebase, our interviews reveal that users need to manually gather specific information at different levels: global level–project background and structural information; local level–key files and component details; and detailed level–information clarification. This suggests that an LLM-powered Codebase Understanding System should support a structured three-level chain of comprehension, reducing users’ manual effort in seeking information across these levels.
@@
Design Opportunity 2: LLM-powered Codebase Understanding System should support information decomposition to align with the human mental model, deepening understanding within each abstraction layer. Our interview findings show that users need to break the codebase structure into different functional components so they can focus on specific parts for deeper understanding. This process is often manual and requires expertise. Therefore, an LLM-powered Codebase Understanding System should automate information decomposition to reduce the manual effort required from users.
@@
Design Opportunity 3: LLM-powered Codebase Understanding System should support information representation connecting across multiple abstraction layers. Information in a codebase is inherently complex, spanning multiple files, functional components, and detailed function implementations. Grasping these relationships requires substantial cognitive effort, particularly when reasoning about inter-module dependencies. Therefore, an LLM-powered Codebase Understanding System should provide structural representations across multiple abstraction layers, enabling easier and more efficient comprehension.
@@
Design Opportunity 4: LLM-powered Codebase Understanding System should maintain an analytical structure and scaffold to reduce manual effort in formulating LLM queries and support transitions between the system’s structured understanding flow and users’ own evolving reasoning. Our interview findings show that querying for specific information often distracts users from their original understanding flow when interacting with conversational LLMs. Moreover, users need both expertise and skill to formulate effective queries–for example, locating the specific information they wish to focus on and translating their analytical goals into language that the LLM can interpret. This challenge is more pronounced for novices, who may not possess the same analytical skills or domain expertise as professionals. To address this, an LLM-powered Codebase Understanding System could provide predefined information structures or commonly requested details that are automatically displayed, allowing users to access relevant insights without manually formulating queries.